18:44:12Q.IF COT(X+Y)=0,THEN THE VALUE OF SIN(X+2Y)?

Grade:12

2 Answers

yash jain
11 years ago

its answer is 1/√2

Shaurya Gupta
11 years ago

1/rt2 is the answer. First cot 90 =0 therefore x+y=pi/2 

open cos(x+y) = 0 

put x=pi/2-y in sin(x+2y) and then same for y. u will get cosy=sinx

solve furthur by opening cos(x+y) identity substitute and u will get 1/rt2